Given K(x)=-5x-6
Find K(3)-K(x)/5
Can someone help me with this...please!

Answer by tutorcecilia(2152)

K(x)=-5x-6
K(3)= -5(3)-6 = -15-6 = -21
K(x)/5 =(-5x-6)/5
.
K(3)-K(x)/5
(-5(3)-6) - (-5x-6/5) [Simplify]
(-15-6) + 5x+6/5
-21+ 5x+6/5
[(-21)(5)/5]+ (5x+6/5)
[(-105)/5]+ (5x+6/5)
(-105+ 5x+6)/5
-99+6/5